OpenStudy (anonymous):

y= 3x^2+6x. I need help finding vertex. Keep getting sign errors using -b/2a formula. Help please! y= 3x^2+6x. I need help finding vertex. Keep getting sign errors using -b/2a formula. Help please! @Mathematics

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

\[\large vertex = (\frac{-b}{2a}, f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

\[\frac{-6}{2(3)} = -1\]

OpenStudy (anonymous):

\[\large f(-1) = -3\]

OpenStudy (anonymous):

\[\large vertex = (-1,-3)\]